The effective length of a magnet is 31.4 cm and its pole strength is 0.5Am. The magnetic moment, when it is bent in the form of a semicircle will be?

Grade:12th Pass

2 Answers

Arun
5 years ago
Dear student
 
Effective length = 2R = 2L/pi
 
Magnetic moment = 2 * 0.314 * 0.5 /3.14 = 0.1 Am2
 
Hope it helps
 
Regards
Arun
Khimraj
5 years ago
Magnetic moment = M = pole strength × effective length = 0.5 × 31.4 × 10–2 = 15.7 × 10–2 Am2 when bent in semicircle, new magnetic moment = M' = (2M/π) ∴ M' = [(2 × 15.7 × 10–2)/(3.14)] = 0.1 Am2